OpenStudy (anonymous):

solve the following equation |2x-7|=9

OpenStudy (anonymous):

x = 8 x = -1

OpenStudy (anonymous):

2x-7=9 x=-1 or x=8

OpenStudy (anonymous):

How do I show work to this? Is there any?

OpenStudy (anonymous):

yes, 2x-7=9 2x=9+7 2x=16 x= 8 because it absolute it can be -9 as well 2x-7=-9 2x=-9+7 2x=-2 x=-1